# Decision Quotient: A Regime-Sensitive Complexity Theory of Exact Relevance Certification

### 16 Mar 2026 v1

@@ 8.3 Low Tensor Rank
Definition 8.6 (Tensor Rank Decomposition).
@@
Remark 8.7.
Separable utility (Definition 8.4) is the special case $R=1$ with $w_{1}=1$, $f_{1}(a)=f(a)$, and $\prod_{i}g_{1i}(s_{i})=g(s)$.
Theorem 8.8 (Low Tensor Rank Reduction).
@@
The artifact certifies three ingredients. First, bounded-rank tensor contraction admits a polynomial bound (L: DQ75.) Second, the low-rank utility representation reduces the relevant optimization step to factored computation in $O(|A|\cdot R\cdot n)$ steps (L: DQ76-77.) Third, these are exactly the paper-specific ingredients needed to invoke the standard tensor-network algorithms cited in the text. Thus the formal support here is a certified reduction-and-transfer statement rather than an end-to-end formalization of the tensor-network backbone itself.  $\blacksquare$
8.4 Tree-Structured Dependencies
@@
Definition 8.9 (Tree-Structured Dependencies).
@@
Theorem 8.10 (Tree Structure Tractability).
@@ 8.5 Bounded Treewidth
Definition 8.11 (Interaction Graph).
@@
Definition 8.12 (Pairwise Utility).
@@
Theorem 8.13 (Bounded Treewidth Reduction).
@@
The artifact certifies the paper-specific reduction from pairwise sufficiency checking to a CSP on the interaction graph (L: DQ80.) It also certifies the transfer of the standard bounded-treewidth complexity bound (L: DQ79, DQ81.) Thus formal support here is a certified reduction-and-transfer theorem: the bounded-treewidth algorithm itself is standard (drawing on backdoor tractability methods [2]), while the paper-specific bridge to that algorithm is mechanized.  $\blacksquare$
8.6 Coordinate Symmetry
@@
Definition 8.14 (Dimensional State Space).
@@
Definition 8.15 (Symmetric Utility).
@@
Definition 8.16 (Orbit Type).
@@
Theorem 8.17 (Orbit Type Characterization).
@@
Theorem 8.18 (Symmetric Sufficiency Reduction).
@@
$\text{orbitType}(s)=\text{orbitType}(s^{\prime})\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$
@@
If $\text{orbitType}(s)=\text{orbitType}(s^{\prime})$, then by Theorem 8.17 there exists $\sigma$ with $\sigma\cdot s=s^{\prime}$. By symmetry, $U(a,s)=U(a,s^{\prime})$ for all $a$, so $\text{Opt}(s)=\text{Opt}(s^{\prime})$.
For sufficiency, we need only check pairs $(s,s^{\prime})$ agreeing on $I$ with different orbit types; same-orbit pairs are automatically equal.  $\blacksquare$
Theorem 8.19 (Symmetric Complexity Bound).
@@ 8.7 Practical Implications
For problems given in the succinct encoding without these structural restrictions, the hardness results of Section 3 apply, justifying heuristic approaches.
9 Implications for Exact Certification
The regime hierarchy matters for more than class labels. The same exact question, which coordinates can be omitted without changing the decision boundary, reappears when one tries to simplify models, build reliable certifiers, check witnesses, approximate exact minimizers, or compress a central interface. This section collects those consequences in one place. The common lesson is that exact relevance is expensive to certify and easy to displace rather than remove. One downstream consequence is a trilemma for exact certifiers in the hard regime: once exact certification is hard, every interface-compression story must be read under the possibility of abstention, weakened guarantees, or overclaiming.
9.1 Exact Simplification and Over-Specification
Proposition 9.1 (Configuration Simplification as Sufficiency-Check).
Let $P=\{p_{1},\ldots,p_{n}\}$ be configuration parameters and let $B$ be a finite set of observable behaviors. Suppose each configuration state determines the subset of behaviors that remain feasible or optimal. Then the question
$\text{``Does parameter subset }I\subseteq P\text{ preserve all decision-relevant behavior?''}$
is an instance of Sufficiency-Check.
(L: CR1, CT12.)
Proof.
Construct a decision problem $\mathcal{D}=(A,X_{1},\ldots,X_{n},U)$ by taking actions $A=B$, coordinate domains $X_{i}$ equal to the domains of the parameters $p_{i}$, and defining $U(b,s)=1$ exactly when behavior $b$ is realized or admissible under configuration state $s$. Then $\text{Opt}(s)$ is the behavior set induced by $s$, and coordinate subset $I$ is sufficient exactly when agreement on the parameters in $I$ forces agreement on the induced optimal-behavior set.  $\blacksquare$
Concrete case study.
Consider a simplified service configuration with three parameters: cache mode $p_{1}\in\{\mathrm{off},\mathrm{on}\}$, retry policy $p_{2}\in\{1,3\}$, and replica count $p_{3}\in\{1,2\}$. Let the observable behavior set be $B=\{\mathrm{latency\text{-}ok},\mathrm{write\text{-}safe},\mathrm{cost\text{-}ok}\}$. Suppose enabling the cache affects only latency, increasing retries affects only write safety, and replica count affects write safety and cost. Then Theorem 9.1 turns the simplification question into an exact sufficiency question on the induced decision problem: which parameter subsets preserve the optimal-behavior set? In this toy instance, if the target simplification must preserve write safety and cost but is indifferent to latency, the exact behavior-preserving core is $\{p_{2},p_{3}\}$, while $p_{1}$ is irrelevant to that target behavior set. The example is small enough to inspect by hand, but it is already an instance of the general Lean-backed reduction from behavior preservation to sufficiency. (L: CR1, CT12.)
9.1.1 Formalized toy POMDP example
Definition 9.2 (One-step POMDP).
A one-step POMDP is a tuple $(S,A,O,p,o,r)$ where $S,A,O$ are finite sets of states, actions, and observations, $p\in\Delta(S)$ is a prior distribution, $o:S\to O$ is the observation map, and $r:A\times S\to\mathbb{R}$ is the immediate reward function. The horizon is one, so the utility of taking action $a$ in true state $s$ is $r(a,s)$.
Write $\text{Opt}_{\mathrm{full}}(s):=\arg\max_{a\in A}r(a,s)$ for the full-information optimal-action set and, for an observation value $o\in O$, set
$\text{Opt}_{\mathrm{obs}}(o):=\arg\max_{a\in A}\mathbb{E}_{S\sim p}[r(a,S)\mid o(S)=o].$
Let $g:O\to O^{\prime}$ be any coarsening of observations and write $\phi:=g\circ o:S\to O^{\prime}$ for the induced coarse signal.
Proposition 9.3 (Reduction to stochastic preservation).
Let $(S,A,O,p,o,r)$ be a one-step POMDP and let $\phi:S\to O^{\prime}$ be as above. Define the decision problem $\mathcal{D}=(A,S,U)$ with $U(a,s):=r(a,s)$ and state distribution $p$. Then $\phi$ is stochastically preserving for $\mathcal{D}$ in the sense of Section 4 if and only if for every $s\in S$ we have
$\text{Opt}_{\mathrm{full}}(s)=\text{Opt}_{\mathrm{obs}}(\phi(s)).$
In particular, checking whether the coarse observation $\phi$ preserves Bayes-optimal actions is exactly an instance of our stochastic preservation predicate for $\mathcal{D}$.
Proof.
By construction $U(a,s)=r(a,s)$ so the full-information optimal-action set in $\mathcal{D}$ equals $\text{Opt}_{\mathrm{full}}(s)$. The coarse-induced optimizer for $\phi$ is, by definition, the set of actions maximizing the conditional expectation of $U$ given the coarse signal. That conditional expectation is exactly the Bayes expectation used to define $\text{Opt}_{\mathrm{obs}}$. Therefore the preservation condition from Section 4 (equality of full-information and coarse-induced optimizers on every state) is equivalent to the displayed identity.  $\blacksquare$
(L: EX1.)
Worked numeric instance.
Consider the one-step POMDP with states $S=\{s_{1},s_{2}\}$, actions $A=\{a,b\}$, prior $p(s_{1})=p(s_{2})=1/2$, observation map $o(s_{1})=o_{1},o(s_{2})=o_{2}$, and reward table
Let $g:O\to O^{\prime}$ be the constant map that coarsens both observations to a single symbol. Then for the coarse signal $\phi=g\circ o$ the conditional expected returns are
$\mathbb{E}[r(a,S)\mid\phi]=\tfrac{1}{2}(2+0)=1,\qquad\mathbb{E}[r(b,S)\mid\phi]=\tfrac{1}{2}(1+3)=2.$
Thus the coarse-induced optimizer is $\{b\}$ while the full-information optimizer at $s_{1}$ is $\{a\}$, so $\phi$ does not preserve Bayes-optimal actions for this instance. This tiny numeric example illustrates how coarse observations can change the optimal-action correspondence and therefore fail the stochastic preservation predicate. (L: EX3.)
This gives a direct, elementary encoding of one-step planning questions as stochastic preservation instances. The explicit-state encoding of the POMDP yields an explicit-state preservation instance, while a succinct encoding of transition or observation structure yields a succinct encoding of the corresponding decision problem, explaining how the same modelling choices generate the different complexity behaviors reported in Section 4.
9.1.2 Formalized reinforcement-learning hyperparameter example
Fix a finite environment set $E$ and finite hyperparameter domains $X_{\alpha},X_{\gamma},X_{\epsilon}$. Let
$H:=X_{\alpha}\times X_{\gamma}\times X_{\epsilon}$
be the set of hyperparameter configurations. For each environment $e\in E$ write $f_{e}:H\to\mathbb{R}$ for the expected return of configuration $h\in H$.
Define the derived decision problem
$\mathcal{D}_{\mathrm{hp}}=(A,H,U),\qquad A:=E,\quad U(e,h):=f_{e}(h).$
Here states are hyperparameter configurations and actions are environments; the utility records return when configuration $h$ is evaluated under environment $e$.
Let $\pi:H\to H^{\prime}:=X_{\alpha}\times X_{\epsilon}$ be the projection that drops the $\gamma$ coordinate. Call $\gamma$ redundant for the tuning task if for every environment $e$ the set of maximizers
$\text{Opt}_{e}:=\arg\max_{h\in H}f_{e}(h)$
projects under $\pi$ to the set of maximizers of the collapsed objective
$g_{e}(h^{\prime}):=\max_{\gamma\in X_{\gamma}}f_{e}(h^{\prime},\gamma),\qquad h^{\prime}\in H^{\prime}.$
Proposition 9.4 (Reduction to a static preservation check).
With notation as above, the statement $\gamma$ is redundant for tuning" is equivalent to a static preservation property for the decision problem $\mathcal{D}_{\mathrm{hp}}$ under the action-space projection $\pi$. Concretely, $\gamma$ is redundant if and only if for every $e\in E$ and every $h\in H$ we have
$h\in\text{Opt}_{e}\quad\Longleftrightarrow\quad\exists h_{0}\in\text{Opt}_{e}\text{ with }\pi(h_{0})=\pi(h),$
which is exactly the condition that the projection $\pi$ preserves the optimal-action correspondence when the decision problem is read with environments as the existential choices.
Proof.
By definition $\gamma$ is redundant when every full maximizer projects to a maximizer of the collapsed objective and conversely every maximizer of the collapsed objective has a witness in the full space; the displayed equivalence is a restatement of that property. The left-to-right implication is immediate from the definition of $g_{e}$, and the right-to-left implication follows because if some projected maximizer has a preimage that attains the collapsed optimum, then there exists a full configuration achieving the same maximum in the original objective. Rewriting these statements as equalities of optimal-action sets for the derived decision problem $\mathcal{D}_{\mathrm{hp}}$ yields the claimed reduction to a static-style preservation check.  $\blacksquare$
(L: EX2.)
The construction above is explicit and elementary. It shows that hyperparameter redundancy questions can be phrased exactly as decision-preservation predicates within our static regime, and that succinct encodings of the environment family or return functions translate into succinct encodings of the corresponding static instance. In practice this lets one import the static-regime hardness and tractability statements directly into the hyperparameter setting.
(L: CR1, CT12.)
Corollary 9.5 (No General-Purpose Exact Configuration Minimizer).
Assuming $\textsf{P}\neq\textsf{coNP}$, there is no polynomial-time general-purpose procedure that takes an arbitrary succinctly encoded configuration problem and always returns a smallest behavior-preserving parameter subset.
(L: DQ36-37.)
Proof.
Such a procedure would solve Minimum-Sufficient-Set in polynomial time for arbitrary succinctly encoded instances, contradicting Theorem 3.5 under the assumption $\textsf{P}\neq\textsf{coNP}$.  $\blacksquare$
Proposition 9.6 (Model-Conditional Rationality of Over-Specification).
Consider a design process in which:
  1. 1.
    removing a parameter requires certifying that it is irrelevant,
  2. 2.
    exact irrelevance certification is charged according to the computational model of Section 2.4, and
  3. 3.
    carrying an extra parameter incurs only linear or otherwise low-order maintenance overhead.
Then, for sufficiently expressive succinctly encoded instances, retaining extra parameters can minimize total cost under the stated cost model.
(L: DQ26, HD14.)
Proof.
By Theorems 3.4 and 3.5, exact sufficiency certification and exact minimization are coNP-complete in the general succinct regime. By Section 7, this hardness is accompanied by exponential lower bounds under ETH in the linear-support regime. If the cost of carrying extra parameters grows only linearly while exact minimization inherits worst-case exponential cost, then beyond a problem-dependent threshold the latter dominates the former. Under that cost model, retaining extra parameters minimizes total cost.  $\blacksquare$
Remark 9.7.
This does not rule out useful simplification tools. It rules out a fully general exact minimizer with worst-case polynomial guarantees. The structured regimes isolated in Section 8 remain viable precisely because they restrict the source of hardness.
9.2 Regime-Limited Exact Certification
Proposition 9.8 (Exact Certification Competence Depends on Regime).
Within the model of this paper, exact certification competence is regime-dependent. In the general succinct regime, exact relevance certification and exact minimization inherit the hardness results of Sections 3 and 7. In the structured regimes of Section 8, exact certification becomes available in polynomial time.
(L: DQ36, DQ40, DQ44.)
Proof.
The negative side is given by Theorems 3.4 and 3.5, together with the ETH-conditioned lower bounds summarized in Section 7. The positive side is given by the tractability results of Section 8, which provide explicit polynomial-time certification procedures under structural restrictions.  $\blacksquare$
9.3 Integrity, Competence, and the Trilemma of Exact Certification
@@
Definition 9.9 (Certifying Solver).
@@
Definition 9.12 (Exact Reliability Claim).
@@
Proposition 9.13 (Integrity and Competence Separation).
@@
Theorem 9.14 (Integrity-Resource Bound (Conditional)).
Fix a regime $\Gamma$ whose exact-sufficiency decision problem is coNP-complete under the declared encoding. Under $\textsf{P}\neq\textsf{coNP}$, no solver is simultaneously integrity-preserving and competent on $\Gamma$ for exact sufficiency.
(L: IC34, DQ31.)
@@
This yields a polynomial-time algorithm for a coNP-complete problem on $\Gamma$, hence coNP $\subseteq$ P. Under $\textsf{P}\neq\textsf{coNP}$, this is impossible. Therefore no solver is simultaneously integrity-preserving and competent on $\Gamma$.  $\blacksquare$
Corollary 9.15 (Exact Reliability Impossibility in the Hard Regime).
@@
(L: DQ28, DQ38.)
Corollary 9.16 (Trilemma of Exact Certification).
@@
(L: IC30, IC32.)
9.4 Witness and Approximation Limits
@@
9.4.1 Witness-Checking Lower Bound
Definition 9.17 (Slot-Inspection Checker).

### 31 Mar 2026 v2

@@ VIII-C Low Tensor Rank
Definition VIII.6 (Tensor Rank Decomposition).
@@
Remark VIII.7.
Separable utility (Definition VIII.4) is the special case $R=1$ with $w_{1}=1$, $f_{1}(a)=f(a)$, and $\prod_{i}g_{1i}(s_{i})=g(s)$.
Theorem VIII.8footnotemark: footnotetext: : DQ75-77 (Low Tensor Rank Reduction).
@@
The artifact certifies three ingredients. First, bounded-rank tensor contraction admits a polynomial bound 606060: DQ75. Second, the low-rank utility representation reduces the relevant optimization step to factored computation in $O(|A|\cdot R\cdot n)$ steps 616161: DQ76-77. Third, these are exactly the paper-specific ingredients needed to invoke the standard tensor-network algorithms cited in the text. Thus the formal support here is a certified reduction-and-transfer statement rather than an end-to-end formalization of the tensor-network backbone itself.  $\blacksquare$
VIII-D Tree-Structured Dependencies
@@
Definition VIII.9 (Tree-Structured Dependencies).
@@
Theorem VIII.10footnotemark: footnotetext: : DQ63, DQ78 (Tree Structure Tractability).
@@ VIII-E Bounded Treewidth
Definition VIII.11 (Interaction Graph).
@@
Definition VIII.12 (Pairwise Utility).
@@
Theorem VIII.13footnotemark: footnotetext: : DQ79-81 (Bounded Treewidth Reduction).
@@
The artifact certifies the paper-specific reduction from pairwise sufficiency checking to a CSP on the interaction graph 646464: DQ80. It also certifies the transfer of the standard bounded-treewidth complexity bound 656565: DQ79, DQ81. Thus formal support here is a certified reduction-and-transfer theorem: the bounded-treewidth algorithm itself is standard (drawing on backdoor tractability methods [2]), while the paper-specific bridge to that algorithm is mechanized.  $\blacksquare$
VIII-F Coordinate Symmetry
@@
Definition VIII.14 (Dimensional State Space).
@@
Definition VIII.15 (Symmetric Utility).
@@
Definition VIII.16 (Orbit Type).
@@
Theorem VIII.17footnotemark: footnotetext: : DQ82 (Orbit Type Characterization).
@@
Theorem VIII.18footnotemark: footnotetext: : DQ83-84 (Symmetric Sufficiency Reduction).
@@
$\text{orbitType}(s)=\text{orbitType}(s^{\prime})\implies\mathrm{Opt}(s)=\mathrm{Opt}(s^{\prime})$
@@
If $\text{orbitType}(s)=\text{orbitType}(s^{\prime})$, then by Theorem VIII.17 there exists $\sigma$ with $\sigma\cdot s=s^{\prime}$. By symmetry, $U(a,s)=U(a,s^{\prime})$ for all $a$, so $\mathrm{Opt}(s)=\mathrm{Opt}(s^{\prime})$.
For sufficiency, we need only check pairs $(s,s^{\prime})$ agreeing on II with different orbit types; same-orbit pairs are automatically equal.  $\blacksquare$
Theorem VIII.19footnotemark: footnotetext: : DQ56, DQ85 (Symmetric Complexity Bound).
@@ VIII-G Practical Implications
For problems given in the succinct encoding without these structural restrictions, the hardness results of Section III apply, justifying heuristic approaches.
IX The Certification Trilemma
@@
IX-A Definitions
Definition IX.1 (Certifying Solver).
@@ IX-A Definitions
Definition IX.4 (Exact Reliability Claim).
@@
Proposition IX.5footnotemark: footnotetext: : DQ10, IC33 (Integrity and Competence Separation).
@@
IX-B Main Impossibility Theorem
Theorem IX.6footnotemark: footnotetext: : IC34, DQ31 (Integrity-Resource Bound (Conditional)).
Fix a regime $\Gamma$ whose exact-sufficiency decision problem is coNP-complete under the declared encoding. Under $\mathrm{P}\neq\mathrm{coNP}$, no solver is simultaneously integrity-preserving and competent on $\Gamma$ for exact sufficiency.
@@ IX-B Main Impossibility Theorem
This yields a polynomial-time algorithm for a coNP-complete problem on $\Gamma$, hence coNP $\subseteq$ P. Under $\mathrm{P}\neq\mathrm{coNP}$, this is impossible. Therefore no solver is simultaneously integrity-preserving and competent on $\Gamma$.  $\blacksquare$
Corollary IX.7footnotemark: footnotetext: : DQ28, DQ38 (Exact Reliability Impossibility in the Hard Regime).
@@
Corollary IX.8footnotemark: footnotetext: : IC30, IC32 (Trilemma of Exact Certification).
@@
IX-C Regime-Dependent Abstention
Proposition IX.9footnotemark: footnotetext: : IC32, DQ2 (Abstention Frontier Expands with Regime Complexity).
$\text{static}\to\text{stochastic}\to\text{sequential}.$
Proof.
As the decision predicate moves from coNP to $P$ to PSPACE hardness, exact worst-case certification requires strictly stronger resources. Fixed polynomial resources therefore force abstention on a larger instance family in higher regimes.  $\blacksquare$
IX-D Connection to Simplicity Tax
The trilemma directly relates to the simplicity tax (Section X-D). When coordinates relevant for exact decisions are omitted from a central model, the unresolved burden must be handled elsewhere: by local resolution, extra queries, or abstention. In the hard exact regime, Theorem IX.6 implies this burden cannot be discharged for free by a polynomial-budget exact certifier. The principle of hardness conservation names this constraint: irreducible certification cost is conserved across the interface, it cannot be eliminated, only displaced.
The trilemma thus operationalizes the theoretical impossibility for practical certifiers: exact relevance certification imposes a structural choice among soundness, completeness, and efficiency, with the precise boundary determined by the regime-sensitive complexity map.
X Structural Consequences for Exact Certification
The regime hierarchy yields both theoretical barriers and practical implications. This section collects structural consequences that follow from the complexity classifications: witness-checking lower bounds, approximation gaps, configuration-simplification limits, regime-dependent certification competence, and the simplicity tax principle. The common theme is that exact relevance certification imposes constraints that cannot be circumvented without either restricting the regime or accepting tradeoffs.
X-A Exact Simplification and Over-Specification
Many practical simplification questions are instances of sufficiency checking. For example, configuration parameter reduction asks whether a subset of parameters preserves all decision-relevant behavior, which directly reduces to Sufficiency-Check (Appendix B contains detailed reductions). This connection yields immediate complexity consequences.
Corollary X.1footnotemark: footnotetext: : DQ36-37 (No General-Purpose Exact Configuration Minimizer).
Assuming $\mathrm{P}\neq\mathrm{coNP}$, there is no polynomial-time general-purpose procedure that takes an arbitrary succinctly encoded configuration problem and always returns a smallest behavior-preserving parameter subset.
Proof.
Such a procedure would solve Minimum-Sufficient-Set in polynomial time for arbitrary succinctly encoded instances, contradicting Theorem III.5 under the assumption $\mathrm{P}\neq\mathrm{coNP}$.  $\blacksquare$
Remark X.2footnotemark: footnotetext: : CR1, CT12, DQ26, HD14.
The corollary does not rule out useful simplification tools; it rules out a fully general exact minimizer with worst-case polynomial guarantees. The structured regimes isolated in Section VIII remain viable precisely because they restrict the source of hardness. Moreover, when certification is computationally expensive but parameter maintenance is cheap, over-specification can be cost-optimal—another manifestation of the simplicity tax principle.
X-B Regime-Limited Exact Certification
Proposition X.3footnotemark: footnotetext: : DQ36, DQ40, DQ44 (Exact Certification Competence Depends on Regime).
Within the model of this paper, exact certification competence is regime-dependent. In the general succinct regime, exact relevance certification and exact minimization inherit the hardness results of Sections III and VII. In the structured regimes of Section VIII, exact certification becomes available in polynomial time.
Proof.
The negative side is given by Theorems III.4 and III.5, together with the ETH-conditioned lower bounds summarized in Section VII. The positive side is given by the tractability results of Section VIII, which provide explicit polynomial-time certification procedures under structural restrictions.  $\blacksquare$
X-C Witness and Approximation Limits
@@
X-C1 Witness-Checking Lower Bound
Definition X.4 (Slot-Inspection Checker).